OpenStudy (anonymous):

A circle has center (a,0), a is greater than 0 and radius 4 units. (a) What is the equation of this circle? (b) Show that if y=x is a tangent to this circle then a=4√2

OpenStudy (anonymous):



OpenStudy (lalaly):

\[(x-a)^2+y^2=16\]

OpenStudy (lalaly):

thats (a)

OpenStudy (anonymous):

can ou show me how you got that?

OpenStudy (lalaly):

if u have a circle with center (h,k) and radius r equation of circle is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

i got as far as the equation is (x-4)^2+y^2=16

OpenStudy (lalaly):

why 4?

OpenStudy (lalaly):

the center is (a,0)

OpenStudy (anonymous):

but the drawing, i guessed a would be 4

OpenStudy (anonymous):

or -4, i honestly guessed though

OpenStudy (lalaly):

no it a x-a

OpenStudy (anonymous):

so what would the answers be?

OpenStudy (anonymous):

and also how?

OpenStudy (anonymous):

lanas right...

OpenStudy (lalaly):

\[(x-a)^2+x^2=1\] try solving this tofind a

OpenStudy (lalaly):

=16

OpenStudy (lalaly):

if u cnt do it ill do it fr u, but everytime i answer the page freezes

OpenStudy (anonymous):

The equation of a circle: \[(x-a)^{2}+(y-b)^{2}=r^{2}\] where (a,b) is your center. In this case (a,0) is your circle. Plug a in for a, and 0 in for b, and the radius in for r \[(x-a)^{2}+(y-0)^{2}=4^{2} = (x-a)^{2}+(y)^{2}=16\]

OpenStudy (lalaly):

http://www.wolframalpha.com/input/?i=%28x-a%29^2%2Bx^2%3D16 frm my equaion the answer is right \[a=4\sqrt2\]

OpenStudy (anonymous):

a=4√2 is a given in this, im looking for x and y, seeing as how y=x is a tangent to the circle